OpenStudy (anonymous):

Slope of (2/3, 4/7) and (2/3, 11/7)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

could you put it up there? this problem is the only thing separating me from my bed x.x

OpenStudy (anonymous):

11/7- 4/7= ? 2/3 - 2/3 =?

OpenStudy (anonymous):

alright alright you would x.x night